OpenStudy (avanti):

How can i find the period of y=tanx without graphing it? I know the answer but is there a way to find it without all the graphing?

OpenStudy (anonymous):

For the general form of tangent, \(y=a\tan b(x-c)+d\), the period is given by \(p=\dfrac{\pi}{b}\).

OpenStudy (avanti):

oh i thought the period is always 2pi/b ? its different for tangent?

OpenStudy (anonymous):

\(\dfrac{2\pi}{b}\) is the period for sine/cosine. Easy to mix up.

OpenStudy (anonymous):

The difference has to do with the singularities of tangent.

OpenStudy (avanti):

oh okay that makes sense now, thanks!

OpenStudy (anonymous):

You're welcome!

OpenStudy (avanti):

@SithsAndGiggles I have another question. How come the period for y=tan(x/3) is 3pi? How to i get rid of the 3 in (x/3) ?

OpenStudy (anonymous):

In this case, \(b=\dfrac{1}{3}\). The period is then \(p=\dfrac{\pi}{\frac{1}{3}}=\dfrac{\pi}{\frac{1}{3}}\cdot\dfrac{3}{3}=\dfrac{3\pi}{1}=3\pi\)

OpenStudy (avanti):

okay got it! thanks so much for all your help!

OpenStudy (anonymous):

You're welcome again!